# Large Language Models Are Strong Audio-Visual Speech Recognition Learners

### v1

@@ Abstract
Multimodal large language models (MLLMs) have recently become a focal point of research due to their formidable multimodal understanding capabilities. For example, in the audio and speech domains, an LLM can be equipped with (automatic) speech recognition (ASR) abilities by just concatenating the audio tokens, computed with an audio encoder, and the text tokens to achieve state-of-the-art results. On the contrary, tasks like visual and audio-visual speech recognition (VSR/AVSR), which also exploit noise-invariant lip movement information, have received little or no attention. To bridge this gap, we propose Llama-AVSR, a new MLLM with strong audio-visual speech recognition capabilities. It leverages pre-trained audio and video encoders to produce modality-specific tokens which, together with the text tokens, are processed by a pre-trained LLM (e.g., Llama3.1-8B) to yield the resulting response in an auto-regressive fashion. Llama-AVSR requires a small number of trainable parameters as only modality-specific projectors and LoRA modules are trained whereas the multi-modal encoders and LLM are kept frozen. We evaluate our proposed approach on LRS3, the largest public AVSR benchmark, and we achieve new state-of-the-art results for the tasks of ASR and AVSR with a WER of 0.81% and 0.77%, respectively. To bolster our results, we investigate the key factors that underpin the effectiveness of Llama-AVSR: the choice of the pre-trained encoders and LLM, the efficient integration of LoRA modules, and the optimal performance-efficiency trade-off obtained via modality-aware compression rates.
@@ I Introduction
Llama-AVSR achieves new state-of-the-art results on the LRS3 dataset for the tasks of ASR (0.81%) and AVSR (0.77%) by training only $42$ and $57$ million parameters, respectively. For VSR, our method outperforms prior works using LLM [36] and attains comparable performance with state-of-the-art methods [3, 4]. Moreover, we experimentally analyze the key factors that lead to the effectiveness of Llama-AVSR for the three tasks, observing that: ❶) the choice of the pre-trained audio and video encoders as well as of the LLM has a big impact on the final performance, ❷) incorporating LoRA modules into the LLM and video encoder are highly beneficial to improve the overall performance whilst requiring limited additional parameters, ❸) the selection of the compression rate is crucial to finding the optimal trade-off between performance and efficiency for all three tasks.
@@ II Method
Modality-specific Projector. The projector, sometimes called connector [24, 29], bridges the pre-trained encoders and the LLM by translating audio and visual features into understandable tokens that the LLM can process. The quality of these tokens significantly influences the MLLM’s performance. Furthermore, the projector plays a crucial role in terms of efficiency since it determines how many tokens will be processed by the LLM, which handles most of the computational load. For instance, $6$ seconds of audio-visual features yield $450$ frames, thus processing these long sequences with the LLM presents substantial computational challenges. Given their simplicity and popularity, we opt to employ linear projectors, which excel at capturing fine-grained details by preserving local audio and visual patterns without loss due to their inherent frame-wise transformation [40]. However, they usually struggle in terms of efficiency and scalability as the number of features (before the projector) and tokens (after) remains constant. We tackle this by downsampling the audio and video features to reduce their sequence length: we first concatenate $K$ (we call it compression rate) consecutive features along the hidden dimension (i.e., the sequence length/hidden size is reduced/increased by a factor $K$), and then the projector maps the audio/video features into audio/video tokens thorough two linear layers to match the text tokens hidden size. Finally, audio and video tokens are concatenated with the textual tokens.
@@ III-B Main Results
auto-avsr [4] Transformer 243 1902/3448 1.0/1.0
@@
Llama-AVSR Whisper 42 1756 0.81
@@
auto-avsr [4] Transformer 250 433 36.3
@@
auto-avsr [4] Transformer 250 1902 23.5
@@
AV-data2vec [11] Transformer 325 433 1.3
auto-avsr [4] Transformer 425 1902/3448 1.0/0.9
@@
We report the results in terms of Word Error Rate (WER) obtained by Llama-AVSR in Table I. We compare our approach with multiple state-of-the-art works based on the considered task (ASR, VSR, AVSR) and the number of labeled hours. We also include the number of trainable parameters and the encoder(s) employed by each method, which may utilize a pre-trained model (as in our case, VSP-LLM [36] and Whisper-Flamingo [46]) or be trained from scratch [4, 2, 8] (i.e., Transformer or Conformer). For the task of ASR (“audio-only setting”), Llama-AVSR sets a new state-of-the-art with a WER of $0.81$% by training on LRS3 + VoxCeleb2 ($1756$ hours), and it also outperforms the other methods when using $433$ and $30$ hours achieving WER results of $1.5$% and $1.1$%, respectively. Remarkably, our approach only requires 42M trainable parameters, which are far fewer compared to all the other methods. We also report the WER achieved by fine-tuning completely a Whisper encoder-decoder model on LRS3 as reported in [46]. Not only does this approach attain a WER much higher than Llama-AVSR ($2.3$% vs $1.1$%), but it also requires the updates of more than $1.5$B parameters. Moreover, their model harnesses Whisper-Large while Llama-AVSR uses the Medium-size version. We also report the case in which AV-HuBERT is used as the audio encoder in place of Whisper (“AV-HuBERT A”).
@@
Finally, for the task of AVSR, we obtain two new SOTA of $0.95$% and $0.77$% when using $433$ and $1756$ hours. This is achieved by using Whisper as the audio encoder and AV-HuBERT as the video encoder. If, instead, we use AV-HuBERT to process both audio and video (“AV-HuBERT AV”), the results are slightly worse, in line with what we observed for the ASR task. We point out that our model outperforms methods that exploit tens of thousands of hours [7, 8]. By comparing with the ASR results, we notice that the additional use of video data is more helpful when we train on LRS3 only ($1.1\%\to 0.95$%), whereas the gain is reduced when more data are available ($0.81\%\to 0.77$%), and this trend in line with previous works [11, 4].
@@ III-C Analysis on the Key Factors of Llama-AVSR
Efficiency-Performance Trade-off for ASR and VSR Tasks. Since most of the computational load lies in the LLM, it is quite common to reduce the number of tokens. At the same time, the loss in resolution inevitably results in a performance’s drop. Consequently, it is crucial to find a compromise in terms of efficiency and performance. In this direction, we report the WER by varying the compression rate $K$ in the range of [1-5] for both the ASR and VSR tasks. For ASR, a compression rate of $1$ (i.e., no compression) means the LLM processes on average $584$ audio tokens, while the LLM processes $117$ tokens for $5$. In Figure 3, we see that $K=3$ seems the best choice in terms of performance and efficiency. Increasing the number of audio tokens causes a modest decrease of the WER for ASR, whereas for VSR we notice a reduction of the WER of around $1$ point. By pushing $K$ to $4$ and $5$, we achieve worse performance for VSR, whereas for ASR we only notice a sharp performance drop when $K$ is set to $5$ for the $40$/$433$h settings. In contrast, the $1756$h setting achieves $1.16$% WER, indicating higher rates are tolerable when more data is utilized.

### v2

@@ Abstract
Multimodal large language models (MLLMs) have recently become a focal point of research due to their formidable multimodal understanding capabilities. For example, in the audio and speech domains, an LLM can be equipped with (automatic) speech recognition (ASR) abilities by just concatenating the audio tokens, computed with an audio encoder, and the text tokens to achieve state-of-the-art results. On the contrary, tasks like visual and audio-visual speech recognition (VSR/AVSR), which also exploit noise-invariant lip movement information, have received little or no attention. To bridge this gap, we propose Llama-AVSR, a new MLLM with strong audio-visual speech recognition capabilities. It leverages pre-trained audio and video encoders to produce modality-specific tokens which, together with the text tokens, are processed by a pre-trained LLM (e.g., Llama3.1-8B) to yield the resulting response in an auto-regressive fashion. Llama-AVSR requires a small number of trainable parameters as only modality-specific projectors and LoRA modules are trained whereas the multi-modal encoders and LLM are kept frozen. We evaluate our proposed approach on LRS3, the largest public AVSR benchmark, and we achieve new state-of-the-art results for the tasks of ASR and AVSR with a WER of 0.79% and 0.77%, respectively. To bolster our results, we investigate the key factors that underpin the effectiveness of Llama-AVSR: the choice of the pre-trained encoders and LLM, the efficient integration of LoRA modules, and the optimal performance-efficiency trade-off obtained via modality-aware compression rates.
@@ I Introduction
Llama-AVSR achieves new state-of-the-art results on the LRS3 dataset for the tasks of ASR (0.79%) and AVSR (0.77%) by training only $42$ and $57$ million parameters, respectively. For VSR, our method outperforms prior works using LLM [36] and attains comparable performance with state-of-the-art methods [3, 4]. Moreover, we experimentally analyze the key factors that lead to the effectiveness of Llama-AVSR for the three tasks, observing that: ❶) the choice of the pre-trained audio and video encoders as well as of the LLM has a big impact on the final performance, ❷) incorporating LoRA modules into the LLM and video encoder are highly beneficial to improve the overall performance whilst requiring limited additional parameters, ❸) the selection of the compression rate is crucial to finding the optimal trade-off between performance and efficiency for all three tasks.
@@ II Method
Modality-specific Projector. The projector, sometimes called connector [24, 29], bridges the pre-trained encoders and the LLM by translating audio and visual features into understandable tokens that the LLM can process. The quality of these tokens significantly influences the MLLM’s performance. Furthermore, the projector plays a crucial role in terms of efficiency since it determines how many tokens will be processed by the LLM, which handles most of the computational load. For instance, $6$ seconds of audio-visual features yield $450$ frames, thus processing these long sequences with the LLM presents substantial computational challenges. Given their simplicity and popularity, we opt to employ linear projectors, which excel at capturing fine-grained details by preserving local audio and visual patterns without loss due to their inherent frame-wise transformation [40]. However, they usually struggle in terms of efficiency and scalability as the number of features (before the projector) and tokens (after) remains constant. We tackle this by downsampling the audio and video features to reduce their sequence length: we first concatenate $K$ (we call it compression rate) consecutive features along the hidden dimension (i.e., the sequence length/hidden size is reduced/increased by a factor $K$), and then the projector maps the audio/video features into audio/video tokens through two linear layers to match the text tokens hidden size. Finally, audio and video tokens are concatenated with the textual tokens.
@@ III-B Main Results
auto-avsr [4] Conformer 243 1902/3448 1.0/1.0
@@
Llama-AVSR Whisper 42 1756 0.79
@@
auto-avsr [4] Conformer 250 433 36.3
@@
auto-avsr [4] Conformer 250 1902 23.5
@@
auto-avsr [4] Conformer 425 1902/3448 1.0/0.9
Fast Conformer [46] Conformer 197 1687 0.9
@@
We report the results in terms of Word Error Rate (WER) obtained by Llama-AVSR in Table I. We compare our approach with multiple state-of-the-art works based on the considered task (ASR, VSR, AVSR) and the number of labeled hours. We also include the number of trainable parameters and the encoder(s) employed by each method, which may utilize a pre-trained model (as in our case, VSP-LLM [36] and Whisper-Flamingo [47]) or be trained from scratch [4, 2, 8] (i.e., Transformer or Conformer). For the task of ASR (“audio-only setting”), Llama-AVSR sets a new state-of-the-art with a WER of $0.79$% by training on LRS3 + VoxCeleb2 ($1756$ hours), and it also outperforms the other methods when using $433$ and $30$ hours achieving WER results of $1.5$% and $1.1$%, respectively. Remarkably, our approach only requires 42M trainable parameters, which are far fewer compared to all the other methods. We also report the WER achieved by fine-tuning completely a Whisper encoder-decoder model on LRS3 as reported in [47]. Not only does this approach attain a WER much higher than Llama-AVSR ($2.3$% vs $1.1$%), but it also requires the updates of more than $1.5$B parameters. Moreover, their model harnesses Whisper-Large while Llama-AVSR uses the Medium-size version. We also report the case in which AV-HuBERT is used as the audio encoder in place of Whisper (“AV-HuBERT A”).
@@
Finally, for the task of AVSR, we obtain two new SOTA of $0.95$% and $0.77$% when using $433$ and $1756$ hours. This is achieved by using Whisper as the audio encoder and AV-HuBERT as the video encoder. If, instead, we use AV-HuBERT to process both audio and video (“AV-HuBERT AV”), the results are slightly worse, in line with what we observed for the ASR task. We point out that our model outperforms methods that exploit tens of thousands of hours [7, 8]. By comparing with the ASR results, we notice that the additional use of video data is more helpful when we train on LRS3 only ($1.1\%\to 0.95$%), whereas the gain is reduced when more data are available ($0.79\%\to 0.77$%), and this trend in line with previous works [11, 4].
@@ III-C Analysis on the Key Factors of Llama-AVSR
Efficiency-Performance Trade-off for ASR and VSR Tasks. Since most of the computational load lies in the LLM, it is quite common to reduce the number of tokens. At the same time, the loss in resolution inevitably results in a performance’s drop. Consequently, it is crucial to find a compromise in terms of efficiency and performance. In this direction, we report the WER by varying the compression rate $K$ in the range of [1-5] for both the ASR and VSR tasks. For ASR, a compression rate of $1$ (i.e., no compression) means the LLM processes on average $584$ audio tokens, while the LLM processes $117$ tokens for $5$. Figure 3 shows that for the ASR task we can compress the audio tokens up to a factor 5 without impacting performance, consistently reducing the number of tokens processed by the LLM. Reasonably, for the 30h setting the pooling factor is more crucial as the model is trained on substantially less data. For VSR, however, increasing $K$ leads to gradually worse results across all settings (the WER increase is around $2.5$-$3$ points when pushing $K$ to $5$), indicating the need for a more careful balance between efficiency and performance.